Circular No. 7443

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 2000cr IN NGC 5395 Marco Migliardi and Alessandro Dimai report their discovery of a su-

pernova (mag ∼ 17) on CCD images taken on June 25.90 and 25.94 UT with the 0.5-m Ullrich telescope at the Col Drusci Observatory in the course of the CROSS program. SN 2000cr is located at α = 13^{h}58^{m}38^{s}.37, δ = +37^{o}26′12′′.4 (equinox 2000.0), which is 2′′ east and 44′′ north of the center of NGC 5395. Nothing was visible at this location on a CROSS image obtained on June 3 (limiting mag 18.5).

W. Li, University of California at Berkeley, reports that SN 2000cr is present (mag ∼ 16.7) on an unfiltered KAIT CCD image obtained on June 26.2 UT. Li provides position end figures for the supernova: 38^{s}.37, 12′′.9, which he notes to be 4′′.2 east and 45′′.2 north of the nucleus of NGC 5395. Li adds that NGC 5395 is interacting with NGC 5394. A KAIT image of the same field taken on June 21.2 showed nothing at the position of SN 2000cr (limiting mag ∼ 19.0).

S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that a spectrum of SN 2000cr, taken by M. Calkins on June 26.17 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-II supernova at an early epoch. The spectrum exhibits a blue continuum with broad P-Cyg H and He features. Narrow superimposed Hα emission yields a recession velocity of 3350 km/s for the host galaxy.

SUPERNOVA 2000cq IN UGC 10354 T. Puckett, Ellijay, GA, reports that he detected SN 2000cq (cf. IAUC

7442) independently on an unfiltered CCD image taken on June 3.127 UT, when it was at mag 19.6, and he provides the following precise position for the supernova: α = 16^{h}21^{m}21^{s}.33, δ = +40^{o}48′09′′.9 (equinox 2000.0).

V1493 AQUILAE R. Novak, Copernicus Observatory, A. Retter, Keele University; and Y.

Lipkin, Wise Observatory, report: “We performed CCD optical photometry of V1493 Aql (N Aql 1999 No. 1) during twelve nights in 1999 July–August. The light curve is modulated with period 0.156 ± 0.001 day. The folded data suggest the presence of eclipses with an amplitude of 0.014 mag.”

2000 June 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green